OpenStudy (anonymous):

In a triangle a cos ^2(c/2)+c cos^(a/2)=3b/2 then its sides a,b,c are in

OpenStudy (anonymous):

@kropot72 ,@Haseeb96

OpenStudy (anonymous):

c^2 = a^2 + b^2 - 2ab cos(C) b^2 = a^2 + c^2 - 2ac cos(B) a^2 = b^2 + c^2 - 2bc cos(A)

ganeshie8 (ganeshie8):

\[\begin{array} \\ a \cos ^2(C/2)+c \cos^2(A/2) &=3b/2 \\ 2a \cos ^2(C/2)+2c \cos^2(A/2) &=3b \\ a (1+\cos C)+c(1+\cos A) &=3b ~~~\color{gray}{\because 2\cos^2\theta /2 = 1+\cos\theta }\\ a + c + (a\cos C + c\cos A )&=3b \\ \end{array}\]

ganeshie8 (ganeshie8):

see if that looks okay so far ^^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b = a cos C +c cos A

OpenStudy (anonymous):

a+c+b=3b

OpenStudy (anonymous):

a+c=2b

OpenStudy (anonymous):

this is in h.p or g.p or a.p form

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

ofcourse it is AP

ganeshie8 (ganeshie8):

if a,b,c are in AP, then 2b = a+c

OpenStudy (anonymous):

yes